Consider the normal distribution of IQs with a mean of 100 and a standard deviation of 16. What percentage of IQs are

a. greater than 95?

b. less than 120

c. between 90 and 110

 



1
Expert's answer
2022-04-10T15:38:14-0400

We have a normal distribution, "\mu=100, \sigma=16."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}."


"a.\ \ z=\cfrac{95-100}{16}=-0.31;\\
P(X>95)=P(Z>-0.31)=\\
=1-P(Z<-0.31)="

"=1-0.3783=0.6217=62.17\%" ​(from z-table).


"b.\ \ z=\cfrac{120-100}{16}=1.25;\\
P(X<120)=P(Z<1.25)=0.8944=89.44\%."


"c.\ \ z_1=\cfrac{90-100}{16}=-0.63;\\
z_2=\cfrac{110-100}{16}=0.63;\\
P(90<X<100)=P(-0.63<Z<0.63)=\\
=P(Z<0.63)-P(Z<-0.63)=\\
=0.7357-0.2643=0.4714=47.14\%."